# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

### v1

@@ 6 Restarted APEX for unknown $\mu^{*}$
Line 11: Terms $\left\|f^{\prime}\left(\bar{y}^{s-1}\right)\right\|^{2}/\left(2\hat{\mu}_{s}\right)$ and $9\mu_{s-1}\Delta_{s-1}/4\hat{\mu}_{s}$ are proper upper bound of $\tilde{\Delta}_{s}$ by the similar proof process for Line 6 if $f$ is $\hat{\mu}_{s}$-QG. Once Algorithm 6 reaches Line 11, by Lemma 6.2, $\mathcal{P}_{s}$ (defined in Line 8) is a $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. From the algorithm procedure, we have $\hat{\mu}_{s}=\mu_{s}/4$. Hence, applying Lemma 4.2 gives a lower bound of $f^{*}$ as
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\Delta_{s},\frac{2}{\hat{\mu}_{s}}\left(\frac{1}{2}\mu_{s}\Delta_{s}\right)\right\}=4\Delta_{s}\ .$ (36)
Line 13: By Lemma 6.2, Algorithm 6 entering Line 13 means that $\mathcal{AGR}$ generates a $\mathcal{\tilde{P}}_{s}$ as a $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. Hence, by Lemma 4.2, if $f$ is $\hat{\mu}_{s}$-QG function, then we have
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\Delta_{s},\frac{2}{\hat{\mu}_{s}}\cdot\frac{1}{2}\mu_{s}\Delta_{s}\right\}\overset{(a)}{=}\Delta_{s}\ ,$ (37)
where $(a)$ holds by $\mu_{s}=\hat{\mu}_{s}$ since once Algorithm 6 enters Line 11, Algorithm 6 will go back to Line 4. ∎∎
@@
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG. In Algorithm 6, if the Half-Flag returned by $\mathcal{AWG}$ or $\mathcal{AGR}$ is False, then we have $\hat{\mu}_{s}>\mu^{*}$. Therefore, for any $s\geq 1$, $\hat{\mu}_{s}\geq\mu^{*}/4$ and the number of times Algorithm 6 updates the QG estimate will not exceed $\left\lceil\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}\right\rceil$.
@@
The proof is based on contradiction. Suppose $f$ is $\hat{\mu}_{s}$-QG, then it follows from Lemma 6.3 that $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$. We will show that, under condition $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$, neither $\mathcal{AWG}$ nor $\mathcal{AGR}$ can return Half-Flag as False. Now, we consider $\mathcal{AWG}$ and $\mathcal{AGR}$ in order.
@@
$\omega_{t}\left(f\left(\hat{x}^{t}\right)-\left(\bar{f}-\left(1+\beta\right)\tilde{\Delta}_{s}\right)\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$ (38)
@@
$\omega_{t}\beta\tilde{\Delta}_{s}-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$ (39)
@@
$\mathcal{AGR}$ is an algorithm that run APEX with level $\tilde{l}=\bar{f}-\theta\tilde{\Delta}_{s}$. Hence, by the similar proof process in Theorem 5.1, we have once $\mathcal{AGR}$ enters Line 10, then we have $\hat{\mu}_{s}$ is overestimated. ∎∎
@@
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG and is $\left(k,L\right)$-PWS. In Algorithm 6, assume that the number of cuts obeys $m\geq k$, and $\hat{\mu}_{1}\geq\mu^{*}$. To find a point set $\mathcal{P}_{s}$ that is an $\left(\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$, Algorithm 6 requires a total number of gradient oracle calls bounded by:
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}+\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\varepsilon^{2}}\right)\ .$ (40)
@@
By Lemma 6.2, when Algorithm 6 enters Line 8 or 13, it produces an $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ . Hence, to obtain an $\left(\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate for point $\bar{y}^{s-1}$, it suffices to enforce $\mu_{s}\Delta_{s}\leq\varepsilon^{2}$.
@@
By Proposition 4.1 with $\iota_{\max}=\sqrt{\left(2\left(1+\gamma\right)\tilde{\Delta}_{s}\right)/\hat{\mu}_{s}}$, together with the third statement in Theorem 5.1 that $\bar{L}\left(t\right)=O(1)L$ if $m\geq k$ , there exist two constants $C_{\gamma},C_{\theta}>0$ such that each call to $\mathcal{AWG}$ or $\mathcal{AGR}$ needs at most $C_{\gamma}m\sqrt{L/\mu}$ or $C_{\theta}m\sqrt{L/\mu}$ inner iterations, respectively, where $\mu$ is the QG estimate used in $\mathcal{AWG}$ and $\mathcal{AGR}$. Let $\bar{C}=\max\left\{C_{\gamma},C_{\theta}\right\}$, then no matter $\mathcal{AWG}$ or $\mathcal{AGR}$, it will return at most $\bar{C}m\sqrt{L/\mu}$ gradient oracle calls.
@@
By Lemma 6.4, the number of quarterings of the QG estimate is bounded by $N_{1}:=\left\lceil\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)\right\rceil$. Let $\tilde{\Delta}^{1,q}$ (resp. $\tilde{\Delta}^{2,q}$) denote $\tilde{\Delta}_{s}$ immediately before (resp. after) the $q$-th quartering of the QG estimates. Since throughout the run we maintain $\hat{\mu}_{s}\geq\mu^{*}/4$, then the total oracle complexity up to achieving $\mu_{s}\Delta_{s}\leq\varepsilon^{2}$ can be bounded by
@@
$\bar{C}m\left(\underbrace{\sqrt{\frac{4L}{\mu^{*}}}N_{1}}_{\text{$\left(\textrm{I}\right)$}}+\underbrace{\sqrt{\frac{L}{\hat{\mu}_{1}}}\log_{1/\theta}\left(\frac{\tilde{\Delta}_{1}}{\tilde{\Delta}^{1,1}}\right)}_{\left(\text{II}\right)}+\underbrace{\sum_{q=1}^{N_{1}-1}\sqrt{\frac{L\cdot 4^{q}}{\hat{\mu}_{1}}}\log_{1/\theta}\left(\frac{\tilde{\Delta}^{2,q}}{\tilde{\Delta}^{1,q+1}}\right)}_{\left(\text{{III}}\right)}+\underbrace{\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\tilde{\Delta}^{2,N_{1}}}{4\varepsilon^{2}}\right)}_{\left(\text{{IV}}\right)}\right).$ (41)
@@
Here $\left(\textrm{I}\right)$ accounts for the quarterings QG estimate; $\left(\text{II}\right)$ is the cost from the start to the first quartering; $\left(\text{{III}}\right)$ accumulates the costs between two quarterings; $\left(\text{{IV}}\right)$ is the cost after the last quartering until $\mu_{s}\Delta_{s}\leq\varepsilon^{2}$.
@@
Let $\bar{\Delta}:=\max_{s\geq 1}\left\{\tilde{\Delta}_{s}\right\}$ be the upper bound of $\tilde{\Delta}_{s}$ in the Algorithm 6 running. Before obtaining $\left(\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate, we have for all $\tilde{\Delta}^{1,q}\geq\varepsilon^{2}/\hat{\mu}_{s}\geq 4\varepsilon^{2}/\mu^{*}$. Moreover, $\sum_{q=1}^{N_{1}-1}\sqrt{\frac{L\cdot 4^{q}}{\hat{\mu}_{1}}}\leq\sqrt{\frac{L\cdot 4^{N_{1}}}{\hat{\mu}_{1}}}\leq\sqrt{\frac{16L}{\mu^{*}}}$. Hence, we have
@@
$\left(\text{II}\right)\leq\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{4\varepsilon^{2}}\right)\ ,\ \left(\text{{III}}\right)\leq\sqrt{\frac{16L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{4\varepsilon^{2}}\right)\ ,\ \left(\text{{IV}}\right)\leq\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{4\varepsilon^{2}}\right).$ (42)
Combining (41) and (42) gives
@@
$S\left(\varepsilon\right)=O(1)\cdot m\left(\sqrt{\frac{L}{\mu^{*}}}\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)+\sqrt{\frac{L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{\varepsilon^{2}}\right)\right)\ .$ (43)
@@
In the following, we give the upper bound for term $\bar{\Delta}$. Algorithm 6 only increases $\tilde{\Delta}_{s}$ at the quartering lines (Line 6 and 11), with update rule
$\tilde{\Delta}_{s}\leq\frac{9\mu_{s-1}\Delta_{s-1}}{4\hat{\mu}_{s}}.$ (44)
@@
Let $\left\{s_{q}\mid 1\leq q\leq N_{1}\right\}$ be the outer iteration index set where quartering occurs, where subscript $q$ means the $q$-th quartering occurs, and write $\tilde{\Delta}_{s_{q}}$ for the value at exit from iteration $s_{q}$. Since $\Delta_{s_{q}-1}\leq\left(1+\beta\right)\tilde{\Delta}_{s_{q}-1}$ and $\hat{\mu}_{s_{q}}=\mu_{s}$, using (44) gives
@@
$\tilde{\Delta}_{s_{q}}\leq\frac{9}{4}\cdot\frac{\mu_{s_{q}-1}}{\mu_{s_{q}}}\Delta_{s_{q}-1}\leq\frac{9\left(1+\beta\right)}{4}\cdot\frac{\mu_{\left(s_{q}\right)-1}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{q}-1}\overset{(a)}{\leq}\frac{9\left(1+\beta\right)}{4}\cdot\frac{\mu_{s_{q-1}}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{q-1}}\ ,$ (45)
@@
$\tilde{\Delta}_{s_{q}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{q-1}\frac{\mu_{s_{1}}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{1}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{N_{1}-1}\frac{4\mu_{s_{1}}}{\mu^{*}}\tilde{\Delta}_{s_{1}}\ .$ (46)
@@
$\tilde{\Delta}_{s_{1}}\leq\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\mu_{s_{1}}}$ (47)
Otherwise, if $s_{1}>1$, by (44) and the initialization bound at stage $s_{1}-1$
$\tilde{\Delta}_{s_{1}}\leq\frac{9\mu_{s_{1}-1}\Delta_{s_{1}-1}}{4\hat{\mu}_{s_{1}}}\overset{(a)}{=}\frac{9\mu_{s_{1}-1}\Delta_{s_{1}-1}}{4\mu_{s_{1}}}\overset{(b)}{\leq}\frac{9\mu_{s_{1}-1}}{4\mu_{s_{1}}}\cdot\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}\overset{(c)}{=}\frac{9}{2\mu_{s_{1}}}\cdot\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}\ ,$ (48)
where $(a)$ holds by $\hat{\mu}_{s_{1}}=\mu_{s_{1}}$ since $\hat{\mu}_{s_{1}}$ is the QG estimate after exiting the $s_{1}$-th outer iteration, $(b)$ follows from $\Delta_{s_{1}-1}\leq\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}$ since the gap always shrinks before the $s_{1}$-th iteration, and $(c)$ holds by $\mu_{s_{1}-1}=\hat{\mu}_{1}$ since before the $s_{1}$-th iteration, the QG estimates do not change. Combining (47), (48) and (46) yields
$\tilde{\Delta}_{s_{q}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{N_{1}-1}\frac{18}{\mu^{*}}\cdot\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}=:\bar{\Delta}\ .$ (49)
Substituting (49) into (43) absorbs the $\beta$ and $\theta$–dependent factors into the hidden $O(1)$, yielding
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)+\log_{1/\theta}\left(\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\varepsilon^{2}}\right)\right)\ .$ (50)
@@
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}+\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\mu^{*}\varepsilon}\right)\ .$ (51)

### v2

@@ 6 Restarted APEX for unknown $\mu^{*}$
Line 11: By Lemma 6.2, Algorithm 6 entering Line 11 means that $\mathcal{AGR}$ generates a $\mathcal{\tilde{P}}_{s}$ as a $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. Hence, by Lemma 4.2, if $f$ is $\hat{\mu}_{s}$-QG function, then we have
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\Delta_{s},\frac{2}{\hat{\mu}_{s}}\cdot\frac{1}{2}\mu_{s}\Delta_{s}\right\}\overset{(a)}{=}\Delta_{s}\ ,$ (6.9)
where $(a)$ holds by $\mu_{s}=\hat{\mu}_{s}$.
@@
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG. In Algorithm 6, if the Quar-Flag returned by $\mathcal{AWG}$ is False, then we have $\hat{\mu}_{s}>\mu^{*}$. Therefore, for any $s\geq 1$, $\hat{\mu}_{s}\geq\mu^{*}/4$ and the number of times Algorithm 6 updates the QG estimate will not exceed $\left\lceil\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}\right\rceil$.
@@
The proof is based on contradiction. Suppose $f$ is $\hat{\mu}_{s}$-QG, then it follows from Lemma 6.3 that $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$. We will show that, under condition $\bar{f}-f^{*}\leq\tilde{\Delta}_{s}$, $\mathcal{AWG}$ can not return Quar-Flag as False.
@@
$\omega_{t}\left(f\left(\hat{x}^{t}\right)-\left(\bar{f}-\left(1+\beta\right)\tilde{\Delta}_{s}\right)\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$ (6.10)
@@
$\omega_{t}\beta\tilde{\Delta}_{s}-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)\leq\bar{L}\left(t\right)\left\|x^{t}-\bar{y}\right\|^{2}\ .$ (6.11)
@@
Proposition 6.1.
Let the weighted sequence $\omega_{t}=(t+2)(t+3)/2$, $\alpha_{t}=4/(t+3)$ and $\theta\in(1/2,1)$ in Algorithm 5. If the objective function $f$ is convex, the Algorithm 5 will return at most $O(1)\left(m\sqrt{\left(6\theta+4\theta\bar{L}(t)/\mu\right)/(2\theta-1)}\right)$ gradient calls in total.
Proof.
Algorithm 5 can be viewed as a specific instance of APEX, where the parameters are set as $\tilde{l}=f\left(\bar{y}\right)-\theta\Delta$, $\alpha_{t}=4/(t+3)$, and $\omega_{t}=(t+2)(t+3)/2$. From Proposition 3.1, we have
$\frac{(t+2)(t+3)}{2}\left[f(\hat{x}^{t})-\left(f(\bar{y})-\theta\Delta\right)\right]-3\theta\Delta\leq\bar{L}(t)\left\|x^{t,0}-\bar{y}\right\|^{2}.$
Before Algorithm 5 terminates, it must hold that $f(\hat{x}^{t})\geq\theta\Delta+\underline{f}=f(\bar{y})-(1-\theta)\Delta$. Therefore,
$\frac{(t+2)(t+3)}{2}\,(2\theta-1)\Delta-3\theta\Delta\leq\bar{L}(t)\left\|x^{t,0}-\bar{y}\right\|^{2}.$
It follows that, as soon as $t\geq\sqrt{\left(4\theta\,\bar{L}(t)/\mu+6\theta\right)/(2\theta-1)}$, we must have $\left\|x^{t,0}-\bar{y}\right\|^{2}\geq 2\theta\Delta/\mu$, establishing the claimed result. ∎
@@
Suppose function $f$ is convex, satisfies $\mu^{*}$-QG and is $\left(k,L\right)$-PWS. In Algorithm 6, assume that the number of cuts obeys $m\geq k$, and $\hat{\mu}_{1}\geq\mu^{*}$. Set $\omega_{t}=(t+2)(t+3)/2,\alpha_{t}=4/(t+3)$ in subroutine $\mathcal{AWG}$ and $\mathcal{AGR}$. To find a point set $\mathcal{P}_{s}$ that is an $\left(2\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$, Algorithm 6 requires a total number of gradient oracle calls bounded by:
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}+\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\varepsilon^{2}}\right)\ .$ (6.12)
@@
By Lemma 6.2, when Algorithm 6 enters Line 8 or 11, it produces an $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ . Hence, to obtain an $\left(2\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate for point $\bar{y}^{s-1}$, it suffices to enforce $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$.
@@
By Proposition 4.1 with $\iota_{\max}=\sqrt{\left(2\left(1+\beta\right)\tilde{\Delta}_{s}\right)/\hat{\mu}_{s}}$, together with the third statement in Theorem 5.1 that $\bar{L}\left(t\right)=O(1)L$ if $m\geq k$ , there exist two constants $C_{\beta},C_{\theta}>0$ such that each call to $\mathcal{AWG}$ or $\mathcal{AGR}$ needs at most $C_{\beta}m\sqrt{L/\mu}$ or $C_{\theta}m\sqrt{L/\mu}$ inner iterations, respectively, where $\mu$ is the QG estimate used in $\mathcal{AWG}$ and $\mathcal{AGR}$. Let $\bar{C}=\max\left\{C_{\beta},C_{\theta}\right\}$, then no matter $\mathcal{AWG}$ or $\mathcal{AGR}$, it will return at most $\bar{C}m\sqrt{L/\mu}$ gradient oracle calls.
@@
By Lemma 6.4, the number of quarterings of the QG estimate is bounded by $N_{1}:=\left\lceil\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)\right\rceil$. Let $\tilde{\Delta}^{1,q}$ (resp. $\tilde{\Delta}^{2,q}$) denote $\tilde{\Delta}_{s}$ immediately before (resp. after) the $q$-th quartering of the QG estimates. Since throughout the run we maintain $\hat{\mu}_{s}\geq\mu^{*}/4$, then the total oracle complexity $S(\varepsilon)$ up to achieving $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$ can be bounded by
@@
$2\bar{C}m\left(\underbrace{\sqrt{\frac{4L}{\mu^{*}}}N_{1}}_{\text{$\left(\textrm{I}\right)$}}+\underbrace{\sqrt{\frac{L}{\hat{\mu}_{1}}}\log_{1/\theta}\left(\frac{\tilde{\Delta}_{1}}{\tilde{\Delta}^{1,1}}\right)}_{\left(\text{II}\right)}+\underbrace{\sum_{q=1}^{N_{1}-1}\sqrt{\frac{L\cdot 4^{q}}{\hat{\mu}_{1}}}\log_{1/\theta}\left(\frac{\tilde{\Delta}^{2,q}}{\tilde{\Delta}^{1,q+1}}\right)}_{\left(\text{{III}}\right)}+\underbrace{\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\tilde{\Delta}^{2,N_{1}}}{8\varepsilon^{2}}\right)}_{\left(\text{{IV}}\right)}\right).$ (6.13)
@@
Here $\left(\textrm{I}\right)$ accounts for the quarterings QG estimate (Quar-Flag returned by $\mathcal{AWG}$ is False and $f(\hat{x}^{s})\leq\tilde{l}$); $\left(\text{II}\right)$ is the cost from the start to the first quartering; $\left(\text{{III}}\right)$ accumulates the costs between two quarterings; $\left(\text{{IV}}\right)$ is the cost after the last quartering until $\mu_{s}\Delta_{s}\leq 2\varepsilon^{2}$; The factor of 2 arises because the number of oracle calls required by $\mathcal{AWG}$ is less than that required by $\mathcal{AGR}$. For clarity and to provide a unified upper bound, we multiply by 2 in our complexity estimate.
@@
Let $\bar{\Delta}:=\max_{s\geq 1}\left\{\tilde{\Delta}_{s}\right\}$ be the upper bound of $\tilde{\Delta}_{s}$ in the Algorithm 6 running. Before obtaining $\left(2\varepsilon/\mu_{s},\varepsilon\right)$-W-certificate, we have for all $\tilde{\Delta}^{1,q}\geq 2\varepsilon^{2}/\hat{\mu}_{s}\geq 8\varepsilon^{2}/\mu^{*}$. Moreover, $\sum_{q=1}^{N_{1}-1}\sqrt{\frac{L\cdot 4^{q}}{\hat{\mu}_{1}}}\leq\sqrt{\frac{L\cdot 4^{N_{1}}}{\hat{\mu}_{1}}}\leq\sqrt{\frac{16L}{\mu^{*}}}$. Hence, we have
@@
$\left(\text{II}\right)\leq\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{8\varepsilon^{2}}\right)\ ,\ \left(\text{{III}}\right)\leq\sqrt{\frac{16L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{8\varepsilon^{2}}\right)\ ,\ \left(\text{{IV}}\right)\leq\sqrt{\frac{4L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{8\varepsilon^{2}}\right).$ (6.14)
Combining (6.13) and (6.14) gives
@@
$S\left(\varepsilon\right)\leq O(1)\cdot m\left(\sqrt{\frac{L}{\mu^{*}}}\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)+\sqrt{\frac{L}{\mu^{*}}}\log_{1/\theta}\left(\frac{\mu^{*}\bar{\Delta}}{\varepsilon^{2}}\right)\right)\ .$ (6.15)
@@
In the following, we give the upper bound for term $\bar{\Delta}$. Algorithm 6 only increases $\tilde{\Delta}_{s}$ at the quartering lines (Line 6), with update rule
$\tilde{\Delta}_{s}\leq\frac{9\mu_{s-1}\Delta_{s-1}}{4\hat{\mu}_{s}}.$ (6.16)
@@
Let $\left\{s_{q}\mid 1\leq q\leq N_{1}\right\}$ be the outer iteration index set where quartering occurs, where subscript $q$ means the $q$-th quartering occurs, and write $\tilde{\Delta}_{s_{q}}$ for the value at exit from iteration $s_{q}$. Since $\Delta_{s_{q}-1}\leq\left(1+\beta\right)\tilde{\Delta}_{s_{q}-1}$ and $\hat{\mu}_{s_{q}}=\mu_{s}$, using (6.16) gives
@@
$\tilde{\Delta}_{s_{q}}\leq\frac{9}{4}\cdot\frac{\mu_{s_{q}-1}}{\mu_{s_{q}}}\Delta_{s_{q}-1}\leq\frac{9\left(1+\beta\right)}{4}\cdot\frac{\mu_{\left(s_{q}\right)-1}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{q}-1}\overset{(a)}{\leq}\frac{9\left(1+\beta\right)}{4}\cdot\frac{\mu_{s_{q-1}}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{q-1}}\ ,$ (6.17)
@@
$\tilde{\Delta}_{s_{q}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{q-1}\frac{\mu_{s_{1}}}{\mu_{s_{q}}}\tilde{\Delta}_{s_{1}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{N_{1}-1}\frac{4\mu_{s_{1}}}{\mu^{*}}\tilde{\Delta}_{s_{1}}\ .$ (6.18)
@@
$\tilde{\Delta}_{s_{1}}\leq\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\mu_{s_{1}}}$ (6.19)
Otherwise, if $s_{1}>1$, by (6.16) and the initialization bound at stage $s_{1}-1$
$\tilde{\Delta}_{s_{1}}\leq\frac{9\mu_{s_{1}-1}\Delta_{s_{1}-1}}{4\hat{\mu}_{s_{1}}}\overset{(a)}{=}\frac{9\mu_{s_{1}-1}\Delta_{s_{1}-1}}{4\mu_{s_{1}}}\overset{(b)}{\leq}\frac{9\mu_{s_{1}-1}}{4\mu_{s_{1}}}\cdot\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}\overset{(c)}{=}\frac{9}{2\mu_{s_{1}}}\cdot\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}\ ,$ (6.20)
where $(a)$ holds by $\hat{\mu}_{s_{1}}=\mu_{s_{1}}$ since $\hat{\mu}_{s_{1}}$ is the QG estimate after exiting the $s_{1}$-th outer iteration, $(b)$ follows from $\Delta_{s_{1}-1}\leq\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}$ since the gap always shrinks before the $s_{1}$-th iteration, and $(c)$ holds by $\mu_{s_{1}-1}=\hat{\mu}_{1}$ since before the $s_{1}$-th iteration, the QG estimates do not change. Combining (6.19), (6.20) and (6.18) yields
$\tilde{\Delta}_{s_{q}}\leq\left(\frac{9\left(1+\beta\right)}{4}\right)^{N_{1}-1}\frac{18}{\mu^{*}}\cdot\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}=:\bar{\Delta}\ .$ (6.21)
Substituting (6.21) into (6.15) absorbs the $\beta$ and $\theta$–dependent factors into the hidden $O(1)$, yielding
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\left(\frac{4\hat{\mu}_{1}}{\mu^{*}}\right)+\log_{1/\theta}\left(\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\varepsilon^{2}}\right)\right)\ .$ (6.22)
@@
$O(1)\cdot m\sqrt{\frac{L}{\mu^{*}}}\left(\log_{4}\frac{4\hat{\mu}_{1}}{\mu^{*}}+\log\frac{\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\mu^{*}\varepsilon}\right)\ .$ (6.23)